A real estate builder wishes to determine how house size is influenced by family income, family size, and education of the head of household. House size is measured in hundreds of square feet, income is measured in thousands of dollars, and education is measured in years. A computer output is shown below.

Regression Statistics

Multiple R

0.865

R Square

0.748

Adjusted R Square

0.726

Standard Error

5.195

Observations

50

ANOVA

df

SS

MS

F

Signif F

Regression

3

3605.7736

901.4434

33.4081

0.0001

Residual

46

1214.2264

26.9828

49

4820.0000

Coeff.

St.Error

t Stat

P-value

Intercept

-1.6335

5.8078

-0.281

0.7798

Family Income

0.4485

0.1137

3.9545

0.0003

Family Size

4.2615

0.8062

5.286

0.0001

Education

-0.6517

0.4319

-1.509

0.1383

What is the predicted house size (in hundreds of square feet) for an individual earning an annual income of 60 (which is measured in thousands of dollars), having a family size of 3, and having 13 years of education?

Select one:

a. 2.42

b. 7.16

c. 48.17

d. 31.22

e. 29.59

At 3% level of significance, which of the following variables can be removed without significantly hurting the model?

Select one:

a. Family income

b. Family size

c. Education

d. Both family income and family size

e. None

Based on the above model, how does increasing family size by 1 member influence the house size, assuming that all other variables remain unchanged?

Select one:

a. House size will increase on average by 426.15 sq. feet

b. House size will increase on average by 5.286 sq. feet

c. House size will change on average by 4.26 sq. feet

d. House size will change on average by 0.865 sq. feet
Answer #1

Let y denotes the house size, x1 denotes family income, x2 denotes family size, x3 denotes education.

The regression equation of y on x1, x2, x3 is:

y = -1.6335 + 0.4485 x1 + 4.2615 x2 - 0.6517 x3.

  • when x1 = 60, x2 = 3, x3 = 13

y = -1.6335 + 0.4485*60 + 4.2615*3 - 0.6517*13 = 29.59 (option e).

  • Since p-value of education is 0.1383 > 0.03, therefore at 3% level it is insignificant. Hence at 3% level of significance, Education (option c) can be removed without significantly hurting the model.
  • Based on the above model, by increasing family size by 1 member house size will change on average by 4.26 sq. feet assuming that all other variables remain unchanged